### v2

# Full private delegated quantum computing tailored from user to industry

Abstract
In this paper, we present a set of private and secure delegated quantum computing protocols and techniques tailored to user-level and industry-level use cases, depending on the computational resources available to the client, the specific privacy needs required, and the type of algorithm. Our protocols are presented at a high level as they are independent of the particular algorithm used for such encryption and decryption processes. Additionally, we propose a method to verify the correct execution of operations by the external server.
1 Introduction
Quantum computing has advanced from a purely theoretical idea to a field with great interest and investment, due to its great properties and applicability to different academic and industrial contexts. However, a major limitation to the enjoyment of such technology is the accessibility of quantum computers, both for individual users and companies, as such devices are extremely expensive even in their smallest versions.
One of the most reasonable proposals to solve this problem is delegated quantum computing. Delegated quantum computing consists of a client with little or no quantum computational resources delegating the quantum operations he needs to perform to an external server with the necessary resources. This would be a quantum analog to the use of classical cloud servers.
However, in several industrial cases it is essential to preserve the privacy of the input data to the calculation, the outputs of the calculation, and even the calculation itself. A clear example could be a company attempting to solve a combinatorial optimization problem, such as the Quadratic knapsack problem, using a Quadratic Unconstrained Binary Optimization (QUBO) formulation [1] adapted to the characteristics of its problem and solved using the Quantum Approximate Optimization Algorithm (QAOA) [2]. In this case, the values of the coefficients would reveal private information about the problem to the external server, potentially related to an economic, telecommunications or transportation project, where this information would be sensitive. This can also become important in certain contexts at user level.
But as we can see, not all operations involve sensitive information. For example, in the QAOA discussed, the information is given in the $RZ$ and $RZZ$ gates, those of the Hamiltonian of the problem. However, no information appears in the $RX$ gates. Moreover, the $RZZ$ gates, when decomposed into two $CNOT$ and one $RZ$ as in the Fig. 1, we can see that the $CNOT$ part does not contain any information either. Therefore, in a QAOA the only part to protect is the $RZ$ part. In general, the gates we are interested in can be efficiently decomposed in this way [3]. In other cases, such as Grover’s algorithm, the information is not contained in rotation angles, but in the circuit structure itself. Because of this, we differentiate between two types of information to be protected in the operations: angle information and structure information. In this way, we can also divide the types of operations into public operations, which can be performed by the external server without receiving private information, and private operations, which contain confidential information.
[…]
Figure 1: $RZZ$ gate decomposition
Another interesting case could be to try to perform distributed quantum computing, to perform a computation using several simultaneous servers, so that only one (or none) has information about the process or the data. This is particularly useful in distributed quantum machine learning [4], where protocols already exist [5, 6].
Classically, the use of homomorphic encryption [7] is usually chosen to encrypt the input data, operate on them and obtain an encrypted output, which can only be retrieved by the client who has the key. To hide the operations, the most effective way is to split the calculation between different servers, so that only part of the operations of the overall process are known to them.
However, in the quantum case we can take advantage of various quantum properties to hide both data and operations by means of quantum properties such as the non-cloning theorem.
Precisely, from these ideas arise two of the main approaches of realizing this private delegated quantum computing [8]. The first is measurement-based blind quantum computation (MBQC) [9, 10], which consists of using measurements and corrections on a graph state to perform a certain quantum computation, and the second is circuit-based blind quantum computation (CBQC) [11, 12, 13], which consists of encrypting the data with a one-time pad as in BB84 [14], to compute and then decrypt it. Encryption can be performed using $X$ and $Z$ gates [13], or even just $Z$ gates [15]. In addition, there is an extension of this second approach, called Full-blind delegating private quantum computation (FDQC) [16], in which it is also possible to hide the operations performed, achieving complete secrecy. It is important to note that in all these methodologies, the client must be able to generate certain quantum states and perform a minimum number of quantum operations, and in some of them being able to store all the qubits of the process. The communication shall be carried out as shown in Fig. 2.
Figure 2: General communication scheme between the client device and the external server device. a) MBQC: the client sends the qubits to the server and the server performs all quantum operations, which are classically controlled. b) CBQC: the client and the server exchange qubits and perform operations according to the client’s instructions. c) Distributed computing: the client controls several servers with a common node, not communicating with each other directly.
Some of these protocols have been successfully demonstrated experimentally [17, 18], both for simple cases and for algorithms such as Grover’s algorithm [19] or Deustch’s algorithm [20]. In addition, they have also been adapted for some academic problems, such as solving Mahalanobis distance [21], and some industrial application problems, such as Variational Quantum Algorithms [22], the QAOA [23] or the Quantum Neural Networks (QNN) [24]. In addition, various equivalences between different protocols have been studied [25].
However, in all these methods very extreme situations are assumed, which can be relaxed in order to get closer to a possible future where the user has quantum devices with certain capabilities. For example, in CBQC cases we have to be able to create the state of the input qubit. But the fact that we can generate that state, in case it is a general one, brings us closer to the idea that we can apply a general one-qubit quantum gate, or at least one general enough to apply to various cases that we will study.
To get closer to such a realistic situation, we will consider a client with a device capable of
  • Generate a qubit in a controlled state.
  • Receive a qubit.
  • Send a qubit.
  • Apply a series of one-qubit (or multi-qubit) gates to the qubits stored in it.
In addition, we will assume that the quantum server is unreliable, so we will want to hide as much information as possible from it for certain given conditions. This server will be able to perform as many quantum operations as we want and will have the capacity to work with as many qubits as we require. In the worst case, if we need more qubits than a server has, we can use several of them connected by a quantum channel to exchange qubits. In addition, we will assume that it has no quantum errors and is able to send and receive as many qubits as it wants on demand. Furthermore, we will assume that the communications take some time to complete, but do not generate quantum errors in the qubits.
The structure of this paper is as follows. First in Sec. 2 we will begin by explaining briefly how the main techniques known in the field work, and then we will propose combinations and modifications of these techniques to adapt them to the needs and requirements of the client, which may be a company in Sec. 3 or a user with fewer resources in Sec. 4. Finally, in Sec. 6 we will give 3 examples of how such protocols would be realized in a particular algorithm.
2 Delegated quantum computing techniques
In this first section we will quickly explain the basics of each delegated calculation technique, so that the reader can keep in mind the various techniques available to perform the subsequent protocols. Each of these techniques can be adapted directly, in whole or in part, to try to carry out the protocols that we will introduce.
2.1 MBQC
Measurement-based quantum computing consists of preparing an entangled quantum state, called a graph state, and performing a sequence of conditional qubit measurements. This method allows performing the same calculations as in the quantum circuit formalism, but performing entanglement only in the state construction. Moreover, since the results of each measurement step are random, we will have to apply a ‘correction’ to subsequent measurements based on the results of the previous ones.
The graph state gets its name from the fact that it can be expressed as a graph, where each node is a qubit and each edge is an entangling operation between the qubits it links. In the graph state we will have the ancilla qubits and the state qubits. The ancilla qubits are the ones that we will measure to form the desired state in the state qubits, since the entanglement will make the modifications in the ancilla qubits translate into modifications in the state qubits. All ancilla qubits are prepared in a uniform superposition, and entanglement is achieved by applying $CZ$ gates.
$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\alpha_{4}$$\alpha_{5}$$\alpha_{6}$$\alpha_{7}$$\alpha_{8}$$\alpha_{9}$$\alpha_{10}$
Figure 3: Generic graph state and graph state for the method [10].
The key to the method is that measuring a uniform superposition state in one direction of the $X-Y$ plane is equivalent to projecting it in that direction. This is similar to applying a gate to the state. This measurement-related gate [26] is the gate
$J(\theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&e^{i\theta}\\ 1&e^{-i\theta}\end{pmatrix},$ (1)
which is achieved by measuring at the $\theta$ angle and correcting the state. Furthermore, we know that any quantum circuit can be constructed from $CZ$ and $J(\theta)$ gates.
Now, we can see that in this case, once the graph state is generated, there is no need to apply more complex operations to it. Moreover, the only multi-qubit operations applied are those we applied at the beginning of the process, and do not contain circuit information by itself.
Given these characteristics, in [10] they study the possibility of using this technique for secure delegated computing. This works as follows: the client sequentially generates the qubits to be sent to an untrusted external server, and the server applies the rest of the operations. However, instead of generating states in uniform superposition, the client generates them randomly from a specific set of possible states to encode the state. In this way, the client can simply communicate to the server which measurements and corrections to perform, making his own decryption. Thus, without having to receive the qubits back, the client can perform the entire calculation, without revealing its state. Furthermore, this protocol can be adapted to the case of distributed quantum computing [27], in which we have several external servers that store and measure the global graph state.
This protocol allows us to verify whether the server is being honest and performing the calculations we ask it to perform [28]. For this purpose, a set of trap qubits will be set, which will give deterministic results known only to the client. If we obtain an unexpected result, we will know that the external server is not performing the designated operations.
However, it requires an enormous amount of resources that can be limiting, because we need a number of ancilla qubits dependent on the depth of the algorithm to be executed.
2.2 CBQC
Circuit-based blind quantum computation techniques are based on the execution of a quantum circuit, so that the quantum state is encrypted with a one-time key. Conceptually, it can be viewed as a realization of the BB84 protocol, where the sender and the receiver of the state are both the client, while the eavesdropper is the external server. In this context, the client can only perform $X$, $Z$ and $SWAP$ operations, in addition to generating arbitrary one-qubit states and being able to store them, so it cannot perform universal quantum computation. Each time the client wants to perform a quantum gate that it cannot apply on the qubits by itself, it will encrypt the qubits to be operated and send them to the external server to be performed. This encryption is performed by applying an $X^{a}Z^{b}$ operation to the qubit to be encrypted, with the pair $(a,b)$ being the random binary key to be stored by the client. After receiving the qubits back, the client will have to decrypt them with the rules presented in [11, 12].
The main limitation of this scheme is that we need to be able to store all the qubits of the circuit. Also, we need a high number of qubit sends and show our operations to the external server, which can be a problem in cases such as the QAOA.
2.3 FBQC
Full-blind delegating private quantum computation allows hiding all the information of the circuit we are executing, both the quantum state itself through the encryption of the one-time pad, as well as the operations [16]. The operations are hidden through the use of 9 qubits, so that each subset of them is associated with a specific operation. In this way, instead of applying an operation, we will apply all the basic operations at the same time. However, only the client will know which qubits are of interest, and therefore, which operation is applied. In fact, part of the key decryption protocol is performed by the external server itself, hiding the decryption as part of the original operations. The main limitation of this technique is the need to multiply the number of qubits to be used, and also the number of qubit sends, as in CBQC.
3 Protocols adapted to industry needs
In this section we are going to analyze the different cases that can occur in the context of a company as a client, with larger quantum resources and problem sizes. First, in Ssec. 3.1 we will study the unconstrained case, only with a memory size problem, which will serve as an initial scheme to understand the basis of the protocols we will introduce. After that, in Ssec. 3.2 we will address the case of having a constraint on the possible client operations. Specifically, the client will not be able to perform multi-qubit operations, in other words, instead of having one quantum computer of many qubits, the client will have many quantum computers of one qubit.
3.1 $M$ qubits total computer
As a simple initial case, we will start by studying the case in which a company wants to perform a certain quantum computation, has a $M$ qubit quantum computer that can perform all available quantum operations (including multi-qubit and measurements), but needs to perform a computation with $N$ qubits, such that $N>M$. In this case, the main limitation faced by this client is not being able to store so many qubits at the same time, so it would need an external quantum memory. This is not as such a case of delegated computing, but a similar need that can be addressed in a similar way, and will allow us to build the basis of all specific protocols for subsequent cases.
To do this, we will only have to use the external quantum computer as memory, so we will send to it the qubits with which we are not operating at that moment, while we will keep in the client’s device the qubits to which we have to perform operations.
However, we want the data embedded in the states we send back to be completely hidden from the external server, so we perform one-time pad encryption like explained in Ssec. 2.2. This way, when we are going to send a qubit to the external server, we will only have to apply an $X^{a}Z^{b}$ operation to the qubit in question with the key $(a,b)$ and then undo it when we receive it back, using the methods of [29, 15, 16, 11]. For the generation of the random keys for encryption, one option is dedicate a qubit to generate the $a$ and $b$. Other one is to generate all the necessary keys prior to the execution of the algorithm and thus be able to perform the protocol with all the qubits. However, this is not possible in all algorithms, such as in an HHL algorithm [30], since we do not know in advance how many repetitions of the circuit we will have to perform, and therefore, the length of the final circuit. The third option is to use an external classical random number generator, which may compromise the randomness of the keys. The fourth option, and best one, will consist of the Protocol 1 every time we have free slots in our computer after sending a round of qubits to the external server.
Protocol 1
Output: Random key for encryption.
  1. 1.
    The client generates a qubit in uniform superposition in each free slot it has and measures it to obtain a list of keys.
  2. 2.
    Repeat step 1 until the client has all the keys he needs.
Also, as our quantum computer has more possible operations, we can encrypt in different ways, including with multi-qubit gates in the case of sending several qubits at once. However, let us restrict them to the $X$ and $Z$ because of the following. Since the external server is also a quantum computer that can perform operations, it will be a good idea to take advantage of it to perform operations in parallel to the client. However, when introducing operations with encrypted states, the decryption protocols become more complicated. We restrict to $X-Z$ encription because it gives us the security we need and the decryption protocols already exist for these gates. However, we cannot simply apply any operation that can be done in parallel, as we have to keep some of them private. In this way, we must choose that the external server only performs the public operations, while the client performs all the private operations.
Therefore, the Protocol 2 will consist of several steps. A first step in which the client iteratively generates $Q$ groups of $M$ qubits in $Q$ rounds, being $Q=\lceil N/M\rceil$, each one being a qubit of the circuit to be executed. Each round, the client applies all the quantum operations of the circuit it can until one of those qubits has to interact with a qubit of another group, at which point the client will send the group of qubits to the server. Before sending these qubits, the client will encrypt them with the one-time pad, applying the encription $X^{a_{i}}Z^{b_{i}}$ to qubit i, so that its key is $(a_{i},b_{i})$. Obviously, we will make the best possible use of those operations that commute with each other, so we will only stop when strictly necessary. The client will repeat this until he has generated $(Q-1)M$ qubits. In each round, if the external server can perform public operations on the qubits it is storing, it can do so, lightening the client’s future operations load.
When the client reachs the last group, of $N-(Q-1)M$ qubits, from the previous group the client will not send to the server the $P$ qubits that are going to have interaction with this last group of qubits, with a maximum $P$ such that the total number of qubits in the client are $M$. After that, the client will generate the qubits of this group and he will ask the server for $QM-N-P$ qubits that it has stored that have to perform operations with any of the qubits that the client has. In this way, we optimize the number of future requests. The client applies all the corresponding operations on these qubits that he has.
The second step will be for the client to determine which are the next qubits he needs to continue with the circuit process and which are the ones he no longer needs from among those he has. After that, the client encrypts the qubits he does not need with the one-time pad and sends them to the server, asking it in exchange for the qubits he does need, and decrypting them. After decrypting them, the client operates with them as far as he can and repeats this second step until the whole circuit is finished. During this whole process, the server performs the public operations it can with the qubits it has.
It is important to emphasize that the whole process of coordination of what operations it can and cannot do is given by the client, since the client only communicates the public operations to be performed when it is the moment when he needs to perform them. In addition, it must be taken into account that these operations, applied on the encrypted qubits, require a different decryption process afterwards, so this can limit the amount and types of operations that the external server can perform.
The third and last step will be to make the necessary measurements of the circuit, for which the client will ask for the groups of qubits he needs and will measure them himself. He can also ask the external server for help, since the qubits are encrypted, so it cannot obtain information about their measurements. The client and the server can start measuring the qubits that are no longer going to be operated after the last associated operation, so that we mix measurement steps with operation steps, thus saving extra requests.
Therefore, the proposal is the following protocol.
Protocol 2
Input: Quantum circuit, private/public operations list, $N$ (qubits needed) and $M$ (qubits of the client).
Output: Result of the circuit execution.
  1. 1.
    For $i$ in range($Q-1$):
    1. (a)
      Generate a group of $M$ qubits, from qubit $iM$ to qubit $(i+1)M-1$.
    2. (b)
      Apply circuit gates until other qubits are needed.
    3. (c)
      Encrypt the qubits applying $X^{a_{j}}Z^{b_{j}}$ to the qubit $j$, with $a_{j},b_{j}\in\{0,1\}$ random keys.
    4. (d)
      Send the group of qubits to the server.
    5. (e)
      (Parallel) Check if the server can execute any public operation with the qubits it has. If so, the server is instructed to perform it.
  2. 2.
    Last round:
    1. (a)
      In the previous round, do not send qubits that may have operations with the ones we are going to generate.
    2. (b)
      Generate the last qubits, from qubit $(Q-1)M$ to qubit $N-1$.
    3. (c)
      If there are $M^{\prime}$ qubits in the client, being $M^{\prime}<M$, the server is asked for $M-M^{\prime}$ qubits that have operations with those of the client. The client decrypts the qubits it receives with the key $(a_{j},b_{j})$ of the corresponding $j$ qubit and the corresponding rules according to the operation that the qubit has undergone.
    4. (d)
      Apply circuit gates until other qubits are needed.
    5. (e)
      (Parallel) Check if the server can execute any public operation with the qubits it has. If so, the server is instructed to perform it.
  3. 3.
    As long as there are still circuit operations to be executed:
    1. (a)
      The client determines which are the next qubits needed to perform operations.
    2. (b)
      The client sends encrypted with a new key the qubits it does not need and requests the qubits it needs from the server.
    3. (c)
      The client decrypts the qubits it receives with the key $(a_{j},b_{j})$ of the corresponding $j$ qubit and the corresponding rules according to the operation that the qubit has undergone.
    4. (d)
      Apply circuit gates until other qubits are needed.
    5. (e)
      If a qubit completes its last circuit operation, it is measured (if it has to be measured). If it is measured by the server, the server communicates the result to the client and the client decrypts it. If it is not necessary to measure it, it will be sent encrypted to the server so as not to take up space on the client.
    6. (f)
      (Parallel) Check if the server can execute any public operation with the qubits it has. If so, the server is instructed to perform it.
We can also observe that, even if we impose that the client does not have the ability to measure the qubits, the server can read those encrypted qubits in such a way that it does not obtain any information. For this reason, in all the cases that we will analyze from this point on, we will not consider the case of not being able to measure as an isolated case, since in all of them the problem is solved in the same way.
Obviously, we would also have the possibility of executing the MBQC method, sending packets of $M$ qubits each time on initialization and delegating all operations to the external quantum computer. However, this presents us with the limitation that we would need a huge number of qubits to perform such operations, which may limit the circuits to be performed. If there is no limitation on the number of qubits, one possibility would be to send the qubits to the external server, which would entangle them and send them back to the client so that the client could measure them.
3.2 $M$ one-qubit total computers
In this case, we will further restrict the capabilities of the client quantum computer, where instead of being an $M$ qubits quantum computer, it will be a set of $M$ single qubit quantum computers without the ability to perform multi-qubit operations.
This will mean that we will have to change Protocol 2 so that all multi-qubit operations are performed by the external server. In advance, this could mean a compromise of privacy in the multiqubit operations. This is due to the parameter containing the gate, for example the angles of the $RZZ$s, or because of the structure of the target operation, for example the shape of a Grover’s oracle. However, we can see that this is not the case from both approaches. The first one is simple to eliminate because, as we commented at the beginning, these gates can be decomposed in a gate sequence such that the angle is isolated in one or several one-qubit gates, which we can execute on the client computer.
The second requires the addition of redundant multiqubit gates in the middle of the circuit. That is, we will add extra gates that all applied together form the identity, for example, two $CNOT$s in a row, and that cannot be distinguished from the original circuit gates. In this way, we will cover the structure underneath our circuits. However, this can be easily overcome by having the server optimizing the circuit it has received, so as to cancel these gate pairs. To avoid this, we will add requests from the server qubits after each of the server operations (trap or private). In this way, the server will not know in which of these requests we have applied an intermediate gate, making it unable to optimize the circuit. An example would be to apply $CNOT$ on the server, request one of the qubits and apply a gate ($RZ$ or Identity), and send it back to apply another $CNOT$ as well. The server will not know if there have been two trap gates or if we have applied a $RZZ$.
However, if we do not have to apply one-qubit gates in most of the circuit it will be impractical to hide the traps in this way. To deal with such cases, we can request such qubits and simply apply a $SWAP$ gate between any two qubits, so that when we resend them, the server never really knows what the original positions of those qubits are, so it will not be able to optimize the circuit. This $SWAP$ gate can always be performed, since it would only consist of exchanging the output ports of the client quantum computers. In order for making impossible to the server to detect which ports are disconnected, for example by sending light pulses or other signals, or by the associated delay times, all the client ports will have to be disconnected and reconnected at the same time. We will consider both these $SWAP$s and these intermediate gates to be private operations.
There is another alternative to all this. This consists of including several auxiliary qubits, whose state will be 0 or 1, to include the trap gates, which in this case will not have to cancel each other. The qubits in 0 will control the trap gates, while the qubits in 1 will control the true gates. However, this method is not safe because of the following. The first thing is that, in case it is done for gates that contain a data in its angle, we will still be giving that data to the server, even if we give it more different ones to mislead. The second thing is that, in case we want to hide the structure, it will not be useful either, since these qubits will always be in an eigenvector of $Z$, so they can be measured without altering the system, allowing the server to measure them without us noticing. In this way, the server can obtain the circuit in case it can somehow determine what these auxiliary qubits are.
Therefore, after making these changes, the protocol would be as follows.
Protocol 3
Input: Quantum circuit, private/public operations list, $N$ (qubits needed) and $M$ (qubits of the client).
Output: Result of the circuit execution.
  1. 1.
    For $i$ in range($Q-1$):
    1. (a)
      Generate a group of $M$ qubits, from qubit $iM$ to qubit $(i+1)M-1$.
    2. (b)
      Apply circuit one-qubit gates until multi-qubit gates are needed.
    3. (c)
      Encrypt the qubits applying $X^{a_{j}}Z^{b_{j}}$ to the qubit $j$, with $a_{j},b_{j}\in\{0,1\}$ random keys.
    4. (d)
      Send the group of qubits to the server.
    5. (e)
      (Parallel) Check if the server can execute any public operation with the qubits it has. If so, the server is instructed to perform it.
  2. 2.
    Last round:
    1. (a)
      Generate the last qubits, from qubit $(Q-1)M$ to qubit $N-1$.
    2. (b)
      If there are $M^{\prime}$ qubits in the client, being $M^{\prime}<M$, the server is asked for $M-M^{\prime}$ qubits that need private operations with those of the client. The client decrypts the qubits it receives with the key $(a_{j},b_{j})$ of the corresponding $j$ qubit and the corresponding rules according to the operation that the qubit has undergone.
    3. (c)
      Apply circuit one-qubit gates until multi-qubit gates are needed.
    4. (d)
      (Parallel) Check if the server can execute any public operation with the qubits it has. If so, the server is instructed to perform it.
  3. 3.
    As long as there are still circuit operations to be executed:
    1. (a)
      The client determines which are the next qubits needed to perform operations.
    2. (b)
      The client performs a random $SWAP$ process the qubits it does not need, sends it encrypted with a new key and requests the qubits it needs from the server.
    3. (c)
      The client decrypts the qubits it receives with the key $(a_{j},b_{j})$ of the corresponding $j$ qubit and the corresponding rules according to the operation that the qubit has undergone.
    4. (d)
      Apply circuit gates until multi-qubit gates are needed.
    5. (e)
      If a qubit completes its last circuit operation, it is measured. If it is measured by the server, the server communicates the result to the client and the client decrypts it.
    6. (f)
      (Parallel) Check if the server can execute any public operation with the qubits it has. If so, the server is instructed to perform it.
As we can see, this protocol protects both client states, through one-time pad encryption, and client operations, through the use of trap gates, intermediate $SWAP$s and local gates in the client device.
4 Protocols adapted to user requirements
Now let us assume a situation in which the client can be a private user. Here, the user will have few quantum computational resources, which will imply having 1 or 2 qubits, either on an own device or by contracting a cloud service to an external company he trusts.
4.1 $2$ qubits total computer
The case of a 2-qubit device that can perform all possible operations with them is a particular case of the problem with $M$ qubits of Sec. 3.1, so the only modification to be made is to impose that $M=2$.
As we mentioned in the first situation, there are several options to deal with the problem of key generation, but in this case, having only 2 qubits, we cannot dedicate one completely to key generation, so the other options take on a different role.
4.2 $2$ one-qubit total computers
This is, again, a particular case of a prior situation. The requirement can be thought as the $M$ one-qubit computer problem, but for $M=2$. No extra considerations are necessary.
4.3 $1$ qubit total computer
This is one of the most extreme cases, since the client would only be able to store one qubit at a time. This already implies that he will not be able to dedicate a qubit exclusively to key generation. Apart from that, the case of $M$ one-qubit computers with $M=1$ so Protocol 3 still applies.
However, a very important problem arises: we cannot use the $SWAP$s method to disguise the structure of the client’s operations. Therefore, in cases where we need to cover up the structure of the operations, we will have to rely only on being able to introduce unknown gates for the server or use an MBQC method, with the cost in server resources that this implies. We can also try to combine the Ssec. 4.5 technique with the client qubit to deal with the $SWAP$s problem by making it an external node.
4.4 $M$ one-qubit partial computers
Let us assume the case of having $M$ one-qubit computers that cannot perform all the operations we want. Suppose they can only perform $X$ and $Z$, which is the minimum for our encryption. In this case, although we can hide the structure of the circuit (as we already indicated with $SWAP$s), we cannot hide a qubit gates angles, since we cannot perform them ourselves. To hide this information, we can use $SWAP$s so that instead of applying $RZ(\theta)$ on a single gate, we apply $RZ(\theta_{1})RZ(\theta_{2})\dots RZ(\theta_{n})$ on several gates, being $\theta=\sum_{i}^{n}\theta_{i}$, performing intermediate $SWAP$s so that the server cannot know which angles to add to reconstruct the original $\theta$. The rest will be the Protocol 3.
However, this cannot be done in the case of having a single qubit, since we will not be able to do $SWAP$, resulting in having to disclose this information.
4.5 00 qubits case
The most extreme case we can find is the one where the client does not have any quantum resource. In this case, we will need two untrusted servers that can communicate through a quantum channel, but not classically. This can be achieved by having them send the qubits to a common node that can go to any other set of servers, so that they do not know which server they are going to or from which server they are receiving the qubits. The common node performs the $SWAP$s of the communication, informing to the client what the new qubit labels are. We will assume that each server has the capacity to perform all calculations. In this case, the following protocol will be 4. All private operations from structure are taken as public because we have added trap gates as explained in the Protocol 3, following the same logics.
Protocol 4
Quantum circuit, private/public operations list and $N$ (qubits needed).
  1. 1.
    Server 1 generates the $N$ qubits and use them to perform all the public operations it can until it needs to perform a private operation.
  2. 2.
    Server 1 applies the private operation with an angle $\theta_{1}$. If it can perform more than one private operation on the set in this way, it performs it.
  3. 3.
    Server 1 sends its qubits to the common node, who applies the $SWAP$s and sends the qubits to server 2.
  4. 4.
    Server 2 applies the private operation with $\theta_{2}$ such that $\theta=\theta_{1}+\theta_{2}$, thus completing the private operation.
  5. 5.
    Server 2 performs steps 1 and 2 with the following parts of the circuit.
  6. 6.
    Server 2 sends its qubits to the common node, who applies the $SWAP$s and sends the qubits to server 1.
  7. 7.
    All of these steps are repeated until the circuit is completed at server $k$.
  8. 8.
    Server $k$ encrypts the qubits with the one-time pad and sends them to server $k\oplus 1$.
  9. 9.
    Server $k\oplus 1$ measures the encrypted qubits and the client decrypts the result.
As we can see, in this method the operations are completely protected, but the states are not protected until they are going to be measured. However, the qubits are combined randomly, by the common node, so they are relatively secure. To improve the level of protection, it is recommended to unable to the server which perform the final encryption to measure the qubits. Another way is that neither of the two servers involved in the calculation has the capacity to measure, delegating the measurement to a third server that will measure the already encrypted state. In addition, the keys for encryption on the $k$ server will be generated on the measurement server or externally, not in the encryption server. Before encryption, it will be necessary to check that the keys follow a distribution that does not allow the measurement server to obtain information. This case, if a larger number of servers are used, could be modified for the realization of federated quantum machine learning.
5 Verification algorithm
An important issue is to verify whether the external server is actually applying the quantum operations we are asking it to perform. There is a possibility that the server, despite not being able to obtain information, is dishonest and simply returns erroneous results.
This is a difficult problem to tackle, because in cases of a large number of qubits, we cannot simply simulate the system classically or use properties of the problem to verify the solution. However, and although there are other verification methods we can address [31], for our protocols the verification process is remarkably simple, due to the use of $SWAP$ gates.
We are going to couple a classically simulatable verifier circuit to the original circuit we want to run, without them being connected to each other. This verifier circuit has to be similar to the original one, so that if we mix the qubits of both circuits, the circuit will be homogeneous. With this, we will execute the corresponding protocol with the unified circuit, mixing in each send the original qubits and the verifiers, so that the external server does not know which are of each group. In this way, it will not be able to deliberately omit only operations in the original circuit. Finally, we will get the results of the original circuit and the verifier. As the verifier circuit is classically simulatable, if the server has omitted any operation of the verifier circuit, the obtained result will not coincide with the simulated one. Obviously, this will mean that Protocol 2 will have to execute the extra step before each qubit is sent, a $SWAP$ is performed between them.
The verification will consist of the following.
  1. 1.
    We prepare the original circuit we want to run with $N$ qubits.
  2. 2.
    We add the verifier circuit of $N^{\prime}$ qubits, which is classically simulatable and structurally similar to the original. This second circuit will not be entangled with any of the $N$ qubits of the original circuit.
  3. 3.
    We run the corresponding protocol for the set of the two circuits, with a total of $N+N^{\prime}$ qubits, performing $SWAP$ on each qubit send.
  4. 4.
    We obtain the qubits results of the verifier circuit.
  5. 5.
    If the verifier circuit results match those of the simulation, that means that the external server has performed the circuit well. If not, the external server is not performing the operations we commanded it to perform.
This method allows us, in case we find a discrepancy between the simulation result and the server result, to determine that the external server is not doing what we indicated. However, there may be the possibility that we cannot detect it, which is the case in which the server only chooses qubits from the original circuit to not perform the operations. Therefore, the verifier circuit would not be altered and, thus, we would not be able to detect this attack. However, in this case, for $N$ original circuit qubits and $N^{\prime}$ verifier circuit qubits, the probability that the server chooses randomly a operation from the original circuit will be $O\left(\frac{1}{1+\frac{N^{\prime}}{N}}\right)$. Therefore, for $n$ attempts, the probability that all attempts choose gates that are from the original circuit will be $O\left(\frac{1}{\left(1+\frac{N^{\prime}}{N}\right)^{n}}\right)$. If we choose a sufficiently high $N^{\prime}$ with respect to $N$, we can greatly reduce the probability that all modified operations are from the original circuit, and therefore, we do not detect the attack.
Due to the constraint that the verifier quantum circuit has to be classically simulatable, instead of creating a single verifier circuit with $N^{\prime}$ qubits, we can create $K$ verifier circuits of $P$ qubits each, so that the total number of verifier qubits is $N^{\prime}=KP$. In this way, each verifier circuit will be fully simulatable, while the ensemble will allow the verification to be performed.
As a numerical example, in the case of performing $10^{3}$ shots, a reasonable amount, we will have a probability of non-detection of $\frac{1}{\left(1+\frac{N^{\prime}}{N}\right)^{10^{3}}}$ in case the server only decides to remove a single gate. If we have as many verifier qubits as original ones, we will have a chance of non-detection of the attack of $\frac{1}{2^{10^{3}}}$. If we use a more modest number of verifier qubits, $N^{\prime}=N/10$, we will have a non-detection probability of $\frac{1}{1.1^{10^{3}}}=4\times 10^{-42}$. Even with $N^{\prime}=N/100$, the probability is still $4.8\times 10^{-5}$. So we see that the method is extremely sensitive to attacks, even when relatively few verifier qubits are used.
In addition, this verification process can be performed while we are running our original circuit, so we will be getting in real time this verification along with the circuit we want to run.
6 Examples with quantum algorithms
In this section we will exemplify how to configure the application of the Protocol 2 and 3 for three completely different quantum algorithms: Grover’s algorithm, the QAOA and a QNN. For simplicity, we will not include the verification method.
6.1 Grover’s algorithm
Grover’s algorithm [19] is a quantum search algorithm, which consists of the consecutive repetition of two operators in a circuit so that we max the probability of measuring the data we are looking for. The two operators that we will have to repeat will be the Grover’s diffusion operator $U_{s}$, in charge of applying an inversion with respect to the mean, and the search oracle $U_{\omega}$, which will mark with a negative sign the state we are looking for. We can see in Fig. 4 the Grover’s circuit that searches for the states |$\ket{101}$ and |$\ket{110}$.
[…]
Figure 4: Grover’s circuit that searches for the states |$\ket{101}$ and |$\ket{110}$.
[…]
Figure 5: Applied Grover’s circuit that searches for the states |$\ket{101}$ and |$\ket{110}$. The blue parts are executed by the client and the red parts by the server. The divisors indicate the qubit exchange and whether it is encrypted ($K$) or decrypted ($K^{-1}$).
In this way, the information that we want to protect in this circuit will be the oracle $U_{\omega}$ that marks the desired state. As this operator is mainly structural, since the information of the problem is not included in continuous angles, it will be necessary to protect this structure. The process for a 2 qubit total computer in this case will be as follows, represented in Fig. 5:
  1. 1.
    The client creates the first two qubits and applies $H$ gates to each one. We also apply the $X$ gate to the second qubit.
  2. 2.
    Encrypts the qubits and sends them.
  3. 3.
    The client creates the third qubit and applies an $H$ gate to it.
  4. 4.
    Encrypts the qubit and send it.
  5. 5.
    The server applies a $CCZ$ gate.
  6. 6.
    The client receives back the last two qubits, decrypts them, applies a $X$ gate to each one, encrypts them and exchanges the two qubits before returning them to the server.
  7. 7.
    The server applies $CZ$ between the first two qubits and sends them to the client.
  8. 8.
    The client does nothing and returns the qubits.
  9. 9.
    The server applies $CZ$ between the last two qubits and sends them to the client.
  10. 10.
    The client receives back the first two qubits and exchanges them before returning them to the server.
  11. 11.
    The server applies two $CZ$ (which eliminate the two recent operations), a $CCZ$ and sends the first two qubits to the client.
  12. 12.
    The client decrypts the qubits, applies a $X$ gate to the first one, encrypts them and send them to the server.
  13. 13.
    The server applies the Grover’s diffusion operator.
  14. 14.
    We repeat the process until we finally measure.
6.2 QAOA
In the case of the QAOA, the circuit also consists of repeating 2 operators until the desired state is obtained. The first operator is a cost Hamiltonian operator $U_{H}$, which for a QUBO problem consists of applying $RZ$ and $RZZ$ gates. The second operator is the mixing operator $U_{X}$, which consists of several $RX$ gates. Thus, the circuit will be one like the one in Fig. 6.
[…]
Figure 6: QAOA circuit with 3 qubits.
In this case, since the circuit structure will always be the same, we will have to protect only the $\theta$ and $\phi$ angles of rotation. To do so, we will make use of the decomposition 1. The process for a 2 qubit total computer in this case will be as follows, represented in Figs. 7 (original) and 8 (optimized):
  1. 1.
    The client creates the first two qubits and applies the corresponding $H$ and $RZ$ gates to them, then encrypts and sends them to the server.
  2. 2.
    The client creates the third qubit and applies the corresponding $H$ and $RZ$ gates to it while the server in parallel applies the $CNOT$ gate to the first two qubits. The client encrypts its qubit, sends it and asks for the second qubit.
  3. 3.
    The client decrypts the third qubit and applies the $RZ$ gate, while the server applies the $CNOT$ gate. After that, the client encrypts, and sends the qubit.
  4. 4.
    The server applies the $CNOT$ gate.
  5. 5.
    The client asks for the first qubit, decrypts it and applies the $RX$ gate, while the server applies the $CNOT$ gate. The client encrypts its qubit, sends it and asks for the third qubit.
  6. 6.
    The client decrypts the third qubit and applies the $RZ$ gate. After that, the client encrypts, and sends the qubit.
  7. 7.
    The server applies the $CNOT$ gate and sends the two last qubits to the client.
  8. 8.
    The client decrypts the qubits and applies the $RX$ gates.
  9. 9.
    The process is repeated for each layer of the algorithm, until finally measure.
[…]
Figure 7: QAOA applied circuit with 3 qubits. The blue parts are executed by the client and the red parts by the server.
[…]
Figure 8: QAOA applied circuit with 3 qubits optimized to take advantage of the parallelization of operations. The blue parts are executed by the client and the red parts by the server.
6.3 QNN
In the case of a standard quantum neural network, we would normally want to protect the input data for each training step. This data is exclusively the initial encoding data of each qubit, so all the rest of the circuit can be executed by the server. However, if we want to hide the weights of the model as well, we can do so by applying the gates of one qubit by the client. The process for a 2 qubit total computer in this case will be as follows, represented in Fig. 9:
  1. 1.
    The client generates 2 qubits and applies $RX$ and $RY$ gates to them to encode the data and the first layer of weights. It encrypts them and sends them to the server.
  2. 2.
    The client generates a third qubit and applies the corresponding $RX$ and $RY$ gates to it, while the server applies the $CNOT$ gate. The client encrypts and sends its qubit, while requesting the first one.
  3. 3.
    The client applies the $RY$ gate to the first qubit, while the server applies the $CNOT$ gate. The client encrypts and sends its qubit, while requesting the last two qubits.
  4. 4.
    The client applies the $RY$ gate to the last two qubits. The client encrypts and sends its qubits.
  5. 5.
    The server applies the last two $CNOT$ gates.
  6. 6.
    We repeat adding layers until we finally measure.
[…]
Figure 9: QNN circuit with 3 qubits.
7 Conclusions
In this work we have seen several protocols and techniques to achieve a private delegated quantum computation adapted to the client’s resources and to the needs of securing both data and quantum operations. One of the possible future lines of research may be to reduce the number of qubit sends while maintaining privacy levels, since these increase the execution time of the circuits and in reality have a cost. Another line of research could be to try to modify the protocol to improve both the complexity of the protocol and to try to adapt it to be fault-tolerant.
Acknowledgment
This work is framed within the Quipucamayocs, the activities of the QuantumQuipu community.